OpenStudy (varaujo19):

Hey guys I need some insight to this calculus problem. The resistance R of a copper wire at temperature T = 30° C is R = 15 Ω. Estimate the resistance at T = 37° C, assuming that dR/dT at T = 30 is 0.13 Ω/(° C). (Round your answer to two decimal places.) I know that I need first to come up with a base model to then find the differential. However....I have no idea what formula to use other than V=IR. Something tells me that I need to incorporate the T value somewhere but I'm at a loss. Please help. Thanks <3

OpenStudy (oswaldmurphy):

Welcome to openstudy! @TheSmartOne should know this one :D

OpenStudy (dumbcow):

Let dR/dT be the slope, assume a linear function for R \[R = \frac{dR}{dT} (T - T_0) + R_0\] plug in initial values \[R = 0.13(37-30) + 15\] \[R = 15.91\]

OpenStudy (varaujo19):

Thank you so much @dumbcow